OpenStudy (anonymous):

What is the direct variation equation if y varies directly with x and y = –5 when x = –15?

OpenStudy (anonymous):

guid me thought this pls no answers

OpenStudy (anonymous):

-5 -15

Directrix (directrix):

Giving direct answers on OpenStudy is against the rules of the site. If anyone posts direct answers to one of your questions, then click on the Report Abuse tab and report the user.

OpenStudy (anonymous):

i dont want answers i need help guieding thourg

Directrix (directrix):

Attached is an image about direct variation. Read more at: http://www.regentsprep.org/regents/math/algebra/AO4/Ldirect.htm

Directrix (directrix):

Start with the direct variation equation: y = k * x where k is the constant of variation. We have to find k using the x and y values that were given in the problem.

OpenStudy (anonymous):

so its y=-3x

Directrix (directrix):

Given: y = -5, = -15 --------------------- y = k*x -5 = k * (-15) --> solve for k

Directrix (directrix):

No.

Directrix (directrix):

-5 = k * (-15) Divide both sides by -15 k = ?

Directrix (directrix):